OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3) I think it is A but i'm just making sure

OpenStudy (anonymous):

yup its A (1,3) Solution: to cancel y and get the value of x: \[(-2x + y-1)-(-4x+y+1) = 2x-2\] \[x=1\] then substitue x=1 to any of the 2 equations \[-2(1)+y=1\] \[y=1+2=3\]